HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

HIGH ROVIBRATIONAL STATES OF HC_{3}N: ANALYSIS OF THE RESONANCE NETWORK INVOLVING THE STRONG 2ν_{5} OVERTONE

BAND

\underline{L. BIZZOCCHI} , M. MELOSSO, C. PUZZARINI, Dipartimento di Chimica “Giacomo Ciamician”, via F. Selmi 2, 40126 Bologna (Italy); F. TAMASSIA, Dipartimento di Chimica Industriale “Toso Montanari”, viale del Risorgimento 4,

40136 Bologna (Italy); O. PIRALI, M.-A. MARTIN DRUMEL, Université Paris-Saclay, CNRS, Institut des Sciences Moléculaires d’Orsay, 91405 Orsay

(France)

Cyanoacetylene (HC_{3}N) is a major astrophysical tracer which manifests its presence in Space via strong rotational and infrared emissions. One of the most conspicuous features of its infrared (IR) spectrum is the 2ν_{5} overtone bands located at ∼1300 cm^{−1} (7.7 μm), and falling near the edge of mid infrared atmospheric window. Despite extensive past investigation of the HC_{3}N rovibrational spectrum [1, 2], this band still lacks an accurate analysis which takes into account the complex resonance network involving the v_{5} = 2, bending state. Here, we report on a comprehensive study of the third polyad of interacting lelvels which includes the states: (v_{4} = 1, v_{7} = 2), v_{6} = v_{7} = 2, (v_{5} = 1, v_{7} = 3), v_{7} = 6, and v_{5} = 2, located around 1300 cm^{−1}. The analysis is performed on an extensive dataset collected at SOLEIL (FAR/mid-IR) and at the University of Bologna (mid-IR and rotational data). Our objective is to retrieve a set of spectroscopic constants which have both clear physical meaning and excellent spectral predictive capability. To this aim, we adopt the ro-vibrational Hamiltonian already used for the analysis of the first and second resonance systems of HC_{3}N (e.g. [2]). It includes explicitly vibrational and ro-vibrational l-type resonance effects in bending overtones/combinations as well as the numerous anharmonic resonances occurring between accidentally nearly-degenerate states. The results will serve to improve the data already included in the 2020 release of the HI-TRAN database.

References

[1] L. Bizzocchi, F. Tamassia, M. Melosso, et al., Astrophys. J. Suppl. S. 233, 11 (2017).

[2] F. Tamassia, L. Bizzocchi M. Melosso, et al., J. Quant. Spectrosc. Rad. Trans. 279, 108044 (2021).

p-number: p116  Submitted on Wed, 14 Jun 2023 09:11:45 +0200